OpenStudy (anonymous):

HELPPPPPPPPPPP <3 I GIVE MEDALS LOG 2 (x) + log 2 (x-1) = log 2 (2x - 3) + 1

OpenStudy (anonymous):

log a +log b =log (axb) this is one of the property in log's

OpenStudy (anonymous):

@gibbs

OpenStudy (gibbs):

@Walkerman15 @Here_to_Help15

OpenStudy (anonymous):

@sammixboo @samson245 @KamiBug @jordanloveangel @joejoe213 @henrry123 @AndrewTheCookie @appleduardo

OpenStudy (anonymous):

@chefducky

OpenStudy (anonymous):

go to mathway .com

OpenStudy (anonymous):

@CausticSyndicalist

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

i have no clue

OpenStudy (anonymous):

thanks :(

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

have you tried mathway.com

OpenStudy (anonymous):

yes, it gives the answer but not the explanation

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

if you apply the log rule log(a) + log(b) = log(ab) to the left hand side, then subtract log2(2x-3) from both sides, then apply log(a) - log(b) = log(a/b) to the left hand side again, I think you are looking at something like: log2 ((x^2-x)/(2x-3)) = 1. Then, I think you can apply logbM =c ---> b^c = M, to rewrite like whole things like: 2^((x^2-x)/(2x-3)) = 2^1. Then since they have the same base, you can just solve what the base is raised to, right? So at the end you are just looking to solve ((x^2-x)/(2x-3)) = 1. You should def. double check my reasoning, but I'll leave the remaining work to you.

OpenStudy (jordanloveangel):

@celin3 i would go with @Ajk